3. Given integers from 1 to 8. If five numbers are taken, they can it is certain that two of them number 9. Prove or deny.
Answer #1

Note 1+8=9 2+7-9 3+6=9 4+ 5 = 9 We take otros arroto dass in write & 1,2, .-. ,8}-{1,83 0{2,730{3,630{4,53 So, Hole are these